# Fear over Friends: Examining the Perceived Influence of Others on Vaccination Decisions

### v1

@@ 3 Results
Table 2: Percentage of responses that fit into each category across vaccination groups. Note, responses could use multiple categories of reasoning. Hence, the percentages will not add up to 100%. For definitions of each category, see Tables 4 and 3.
Lastly, to capture reasoning that may not have been addressed by our ordinal-scaled questions, we ask an open ended question to all participants: “In your own words, why did you (not) get a COVID-19 vaccine?”. As discussed in Section 2, we take an open-coding approach to group responses into categories. The final set of categories and corresponding percentage of responses in the category can be found in Table 2. Detailed definitions of each category can be found in Tables 4 and 3.
@@
Other less frequent reasons by vaccinated participants included trust in science (14.81% of responses), desire to return to normal (7.41% of responses), and getting vaccinated being “the right thing to do” (5.85% of responses). Notably, participants explicitly stating they got vaccinated due to trusting in their social circle, either through direct advice from social circle, a request to get vaccinated from social circle, or family members also getting vaccinated, only was present in 1.75% of responses. Further, social pressure, both directly from social circles or indirectly from social norms, was mentioned in only 2.73% of responses. Hence, while perceived social influence was high amongst vaccinated participants, it may not be the primary reasoning behind the decision.
@@ 4 Discussion
In this study, we found that vaccinated participants perceived more influence from and agreement with others than unvaccinated participants. The vaccinated population also perceived COVID-19 as more dangerous to others than the unvaccinated population did. For context, we approximated that participants in the vaccinated population had larger social circles on average and lived in more densely populated areas. Hence, vaccinated participants had a higher likelihood of interacting with someone who had been severely ill or died from COVID-19 than unvaccinated participants. These results expand upon and add robustness to the findings of prior work on social influence during pandemic and epidemics tunccgencc2021social .
In contrast, social influence does not seem significant in deciding not to be vaccinated. In open-ended responses, unvaccinated participants most frequently cited fear of the vaccine and distrust in the institutions who make and support the vaccines.
@@
While the primary goal of this study was to examine social influence, it is clear that social influence is not the only mechanism of impression in decision making. Other factors not directly measured in this study that likely also play a role in vaccination decisions include media consumption schulhofer2013newspapers ; hayes2015local ; gentzkow2011effect ; joseph2022local and individual, generational, and cultural memory loftus2005planting ; lesthaeghe2014second ; ruck2020cultural . Our qualitative results support this notion.
Notably, despite vaccinated participants citing social influence at varying levels in their open-ended responses, they too were motivated by fear, frequently citing that they got vaccinated to protect themselves or to avoid dying from COVID-19.
@@
Fear and anxiety being a driver of health-related decision making is well-supported in the literature turk2006using ; devlin2007comparative ; krishen2015fear , including behaviors during the COVID-19 pandemic. For example, Harper et al. (2021) found that from a sample of 324 people from the U.K. that “the only predictor of positive behavior change (e.g., social distancing, improved hand hygiene) was fear of COVID-19 harper2021functional ”, including when controlling for political and ideological variables. When fear and anxiety motivates vaccination, this fear is functional (as Harper et al. (2021) puts it), but, as our study suggests, when fear and anxiety motivates not getting vaccinated, the fear may be counterproductive and even detrimental.
@@
In conclusion, we found that vaccinated populations perceived more influence from their social circles than unvaccinated populations. This finding held true across various social groups, including family, close friends, co-workers, and neighbors. Similarly, we found that vaccinated participants perceived that others agreed with their decision to get vaccinated significantly more than unvaccinated participants perceived that others agreed with their decision to not get vaccinated. Indirect measures of social influence also followed this trend. Vaccinated participants perceived COVID-19 as more dangerous to their social circles than unvaccinated participants. Despite the clear differences in perceived social influence across the groups, we found through open-ended responses that both vaccinated and unvaccinated participants frequently cited fear as a motivating factor in their decision: vaccinated participants feared COVID-19, while unvaccinated participants feared the vaccine itself. Together, our results expand the current literature on vaccination behavior and add robustness to several previous findings.
@@ 5 Limitations
Supplementary information Additional figures are provided in the Supplemental materials.

### v2

@@ 3.6 Additional Context: Open-ended Reasoning
Table 5: Percentage of responses that fit into each category across vaccination groups. Note, responses could use multiple categories of reasoning. Hence, the percentages will not add up to 100%. For definitions of each category, see Tables 6 and 7 at the end of the paper.
To add more context our results and capture reasoning that may not have been addressed by our research questions, we ask an open ended question to all participants: “In your own words, why did you (not) get a COVID-19 vaccine?”. As discussed in Section 2.2, we take an open-coding approach to group responses into categories. The final set of categories and corresponding percentage of responses in the category can be found in Table 5. Detailed definitions of each category can be found in Tables 6 and 7 at the end of this paper.
@@
Other less frequent reasons by vaccinated participants included trust in science (14.81% of responses), desire to return to normal (7.41% of responses), and getting vaccinated being “the right thing to do” (5.85% of responses). Notably, participants explicitly stating they got vaccinated due to trusting in their social circle, either through direct advice from social circle, a request to get vaccinated from social circle, or family members also getting vaccinated, only was present in 1.75% of responses. Further, social pressure, both directly from social circles or indirectly from social norms, was mentioned in only 2.73% of responses. Hence, while perceived social influence and agreement was high amongst vaccinated participants, it was very likely not the primary reasoning behind the decision (as already suggested by Figures 2 and 3).
@@ 4 Discussion
In this study, we found that vaccinated participants perceived more influence from and agreement with others than unvaccinated participants, particularly from strong social ties like one’s family and close friends. These findings broadly align with the findings of previous work. Tunçgenç et al. (2021) showed that the perceived approval and adherence of others to social distancing during COVID-19 predicted participant’s adherence to social distancing, particularly when others within one’s close social circle approved and adhered to pandemic guidelines tunccgencc2021social . However, this study did not demonstrate differences in perceived social influence between groups, rather the authors demonstrated that there is a relationship between individual adherence and perceptions of social circle adherence. We add nuance to this finding, by showing perceptions of influence and approval may not be symmetrically across groups (those who follow pandemic guidelines and those who do not). Our finding of vaccination influences primarily lying within the family is consistent with a study of HPV vaccinations among adolescent girls, whose strongest influence was the mother Karafillakis2022ITT .
@@
In contrast, social influence and agreement does not seem to be the primary factor in deciding not to be vaccinated. In open-ended responses, unvaccinated participants most frequently cited fear of the vaccine and distrust in the institutions who make and support the vaccines. While the primary goal of this study was to examine social influence, it is clear that social influence is not the only mechanism of impression in decision making. Other factors not directly measured in this study that likely also play a role in vaccination decisions include media consumption schulhofer2013newspapers ; hayes2015local ; gentzkow2011effect ; joseph2022local and individual, generational, and cultural memory loftus2005planting ; lesthaeghe2014second ; ruck2020cultural ; horne2023generational . Our qualitative results support this notion.
@@
Notably, despite vaccinated participants citing social influence at varying levels in their open-ended responses, they too were motivated by fear, frequently citing that they got vaccinated to protect themselves or to avoid dying from COVID-19. Fear and anxiety being a driver of health-related decision making is well-supported in the literature turk2006using ; devlin2007comparative ; krishen2015fear , including behaviors during the COVID-19 pandemic. For example, Harper et al. (2021) found that from a sample of 324 people from the U.K. that “the only predictor of positive behavior change (e.g., social distancing, improved hand hygiene) was fear of COVID-19 harper2021functional ”, including when controlling for political and ideological variables. When fear and anxiety motivates vaccination, this fear is functional (as Harper et al. (2021) puts it), but, as our study suggests, when fear and anxiety motivates not getting vaccinated, the fear may be counterproductive and even detrimental.
@@
In conclusion, we found that vaccinated populations perceived more influence from their social circles than unvaccinated populations. This finding held true across various social groups, including family, close friends, and neighbors. Similarly, we found that vaccinated participants perceived that others agreed with their decision to get vaccinated significantly more than unvaccinated participants perceived that others agreed with their decision to not get vaccinated. This finding held across family, close friends, coworkers and neighbors. Indirect measures of social influence did not follow this trend. Vaccinated participants did not perceive COVID-19 as any more dangerous to their social circles than unvaccinated participants. Despite the clear differences in perceived social influence and agreement across the groups, we found through open-ended responses that both vaccinated and unvaccinated participants frequently cited fear as a motivating factor in their decision: vaccinated participants feared COVID-19, while unvaccinated participants feared the vaccine itself. Together, our results expand the current literature on vaccination behavior and add robustness to several previous findings.
@@ 5 Limitations
Third, our open-ended results suggest fear as a major factor in both getting vaccinated and not getting vaccinated. However, given the study was not designed to measure fear rather social influence and agreement, we cannot robustly establish this result outside of our qualitative analysis.
@@
Supplemental Materials: Additional survey description and analysis can be found in the supplement.